### v1

# Privacy Preservation Techniques (PPTs) in IoT Systems: A Scoping Review and Future Directions Author’s draft for soliciting feedback - February 20, 2025

Abstract
Privacy preservation in Internet of Things (IoT) systems requires the use of privacy-enhancing technologies (PETs) built from innovative technologies such as cryptography and artificial intelligence (AI) to create techniques called privacy preservation techniques (PPTs). These PPTs achieve various privacy goals and address different privacy concerns by mitigating potential privacy threats within IoT systems. This study carried out a scoping review of different types of PPTs used in previous research works on IoT systems between 2010 and early 2023 to further explore the advantages of privacy preservation in these systems. This scoping review looks at privacy goals, possible technologies used for building PET, the integration of PPTs into the computing layer of the IoT architecture, different IoT applications in which PPTs are deployed, and the different privacy types addressed by these techniques within IoT systems. Key findings, such as the prominent privacy goal and privacy type in IoT, are discussed in this survey, along with identified research gaps that could inform future endeavors in privacy research and benefit the privacy research community and other stakeholders in IoT systems.
Index Terms:
Scoping, Privacy threats, Privacy-Enhancing Technology (PET), Privacy Preservation Technique (PPT), Privacy Engineering (PE), and IoT.
I Introduction
As the Internet of Things (IoT) continues to integrate into human daily life, effective privacy preservation techniques (PPTs) have become essential [1]. These PPTs mitigate privacy risks by protecting sensitive personal information within IoT systems.
The Internet of Things (IoT) concept dates back to the 1970s, when it evolved from a closed proprietary telemetry system, as explained by McKelvey et al. [2]. However, significant attention to privacy preservation within IoT began to grow after 1999, when the term ”Internet of Things” was first coined by Kevin Ashton [2].
This increase in privacy focus is evident in the growing number of related publications on privacy in general and IoT systems from 1960 to early 2023, as shown in Figure 1. Figure 1 was generated from an analysis of privacy-related publications collected from Google Scholar using Harzing’s Publish or Perish software, designed to retrieve and analyze academic citations[3].
Figure 1: The line graph shows publication trends for Privacy Preservation Techniques. The blue line indicates general privacy publications, while the green line shows privacy-related publications in IoT (Internet of Things) systems.
From the analysis in this study, trends in IoT-related privacy publications started in 2002, as seen in Figure 1, with a significant increase from 2008 to 2023. This increase highlights increased awareness and recognition of the privacy risks associated with IoT systems.
It is important to note that, as illustrated in Figure 1, the collection of data was concluded in 2023. This excluded any recent publications from the second quarter of 2023 until now, as evidenced by the declining slope at the end of the line graph.
As mentioned in some literature, technological advancements contribute to the new privacy challenges by exacerbating privacy risks [4]. Threat actors can now launch a privacy threat leveraging these new technologies such as Generative AI [4, 5]. However, technological advancements have also facilitated the development of sophisticated PPTs [6, 7]. Some of these technologies, such as artificial intelligence (AI), post-quantum cryptography, and blockchain, have been leveraged to build privacy-enhancing technologies (PETs) used to create PPTs that mitigate privacy threats.
This study leverages scoping review, a systematic method for mapping existing evidence on a topic, by identifying key concepts, theories, and knowledge gaps [8]. This study embraces this approach due to its comprehensive overview of a subject area, helping researchers and practitioners to understand current knowledge and highlighting areas for further investigation.
This study aims to use a scoping review to analyze various PPTs in the context of IoT, focusing on the technologies used to build PET, PPT implementation in the computing layers of IoT architecture, privacy protection types, different IoT applications in which PPTs are implemented, and privacy goals of these PPTs. This study addresses gaps not covered by previous reviews on PPT in IoT systems.
Contributions This review will focus on the PPTs related to IoT from the literature. At the time this study was conducted, and to the best of our knowledge, we were unaware of any publication explicitly focusing on a scoping review of PPTs used in IoT applications. Our contribution will focus on
  • providing a scoping review of state-of-the-art PPTs in IoT systems from paper publications in the past thirteen and a half (13.5) years.
  • analyzing studies on PPTs in the literature based on IoT application area, PPT goals, technologies used to build PETs, and the privacy types protected by these PPTs as gathered from the literature.
  • proposing a clear relationship among PPT, PET, privacy threat, privacy goals, and privacy types.
  • presenting the importance of device identity privacy in IoT.
  • comparison of each PPT from the literature to identify the most frequently used and least used PPTs across different IoT applications.
This paper is organized as follows: Section I explains the trends in general privacy publications and those specific to privacy in IoT systems. Section II presents a brief background on privacy and key privacy terms in IoT systems, Section III shows the uniqueness of this study by comparing it to previous reviews on PPTs. Section IV explains the scoping review guideline and framework used in this study. Section V identifies the research questions in this study, while Section VI explains how we gathered relevant studies in this scoping review. Section VII presents the databases and the year range of selection of papers reviewed in this study. Section VIII categorizes the collected papers into five key areas our review focused on: PPT goals, technologies used to build PET, IoT applications and their PPTs, IoT computing layer in which PPTs are implemented, and privacy types which these PPTs. Section IX summarizes our findings, providing answers to the research questions in our study. Section X discusses the future direction based on the research questions in our study. Lastly, Section XI is the conclusion of the study, summarizing the outcome of the study and the limitations of this study.
II Definitions of key concepts
II-A Privacy
Privacy has evolved significantly since its origins in the 14th through 18th centuries. Initially focused on protecting one’s body and home, it transitioned by the end of the 19th century to include the control and protection of personal information [9]. Privacy refers to an individual’s ability to determine how much of their thoughts and actions they share with others and under what circumstances. This concept is closely tied to personal identity and is viewed as safeguarding an individual’s autonomy and independence [10].
II-B Types of Privacy Protection
Privacy in the IoT revolves around users and other entities interacting within IoT systems. Finn et al. [11] identify seven types of privacy in technology. These types include the privacy of the person, the privacy of personal behavior, the privacy of personal communication, the privacy of personal data, the privacy of location and space, the privacy of thoughts and feelings, and the privacy of associations. However, not all of these privacy types are relevant to IoT. Therefore, based on existing literature, this study outlined the specific types of privacy in IoT systems.
  • User privacy: To protect user data by the service provider and provide transparency on the kind of users’ information they are collecting [12].
  • Device Identity privacy: To prevent unauthorized device tracking and smart device information disclosure such as firmware version, operating system, MAC address, etc. [12].
  • Communication channels privacy: To protect communication link conveying information between device-to-device, device-to-cloud, device-to-human and device-to-application [11].
  • Location privacy: To protect the user and smart device location information. This could prevent threats such as profiling [11, 12].
  • Group privacy: To protect the privacy of a group of IoT system users [11]. For example, a family member of a Smart Home (SH) owner or department member in an Industrial IoT (IIoT) that has access to operate an automated system in an industry.
II-C Privacy Threats in IoT
Depending on the threat actor, some privacy threats can be malicious, while some can be non-malicious. Privacy threats such as the identification of the subject, linkage, data leakage, inventory attacks, eavesdropping, life cycle transition, privacy violation during the presentation, and data tampering could be malicious due to the threat actor involved [13, 14, 15, 16]. In contrast, privacy threats such as Localization and tracking, and profiling are not usually malicious when they involve actors such as service providers, security agents, and third-party providers [13, 14].
II-D Privacy Preservation Techniques (PPT)
PPTs are methods or processes implemented in IoT to achieve privacy goals [17]. These privacy goals vary depending on the techniques and the IoT application involved. These privacy goals include data masking, pseudonymization, data perturbation, multipath routing, and access controls, as shown in Table III. To achieve this goal by PPTs, the right technological building block needs to be developed, which are referred to as privacy enhancing technologies (PET) [18, 19, 20, 21]. PET in IoT may be challenging due to IoT systems’ constraints, such as memory capacity, processing power, etc. However, implementing these PETs in a lightweight manner will make them suitable for IoT systems [17].
II-E Privacy Enhancing Technologies (PET)
These are technologies used to achieve PPTs in an IoT system. In IoT, PPTs have evolved, employing diverse PETs (such as differential privacy, Homomorphic encryption, pseudonymization, and secure multi-party computation, to mention a few) to provide different privacy preservation solutions [22, 23, 24, 25, 26, 27, 28].
II-F General Technology used in achieving Privacy Preservation Approach in IoT
The Internet of Things (IoT) utilizes various privacy preservation approaches to achieve specific privacy goals and types. These approaches often leverage PETs developed through technological methods like blockchain [29], machine learning/deep learning [30], cryptography [31], informed consent [32], quantum-safe algorithms [33], mathematical models [34], and privacy-based infrastructure.
III Comparison with existing surveys and reviews
This section compares previous related surveys, emphasizing key parameters such as the technology used to build PET, goals of PPT, specific IoT applications, and the implementation of privacy and PPT across various IoT computing layers.
It is important to provide a brief background on the types of literature reviews and which literature review is used in our study.
Types of literature review: There are 14 types of literature reviews as explained by Grant et al. [35], which can be categorized into four main types: narrative, rapid, scoping, and systematic. A narrative review covers a small selected part of a study to support empirical research. However, replicating it is often challenging and may yield biased results due to its non-comprehensive nature[35, 36]. Rapid reviews quickly summarize and synthesize research findings within the constraints of time and resources, systematically striving for comprehensiveness and details [35, 37]. Systematic reviews aim to identify, appraise, and synthesize all empirical evidence that meets specific inclusion criteria to address focused research questions [35, 36]. Scoping reviews employ a comprehensive search strategy to map the literature within a broad context but do not synthesize evidence, often serving as a preliminary step before conducting a more focused systematic review [35, 38]. This study employs a scoping review methodology to map and analyze existing literature on PPTs within IoT systems.
Paper PPT Review PPT Goals IoT applications Privacy Types PPT implementation in IoT computing layers Literature Review
[18] $\star$ $\star$ Systematic Mapping
[19] $\star$ $\star$ $\star$ Narrative Review
[20] $\star$ Systematic Mapping
[21] $\star$ Systematic Review
Our work ① ③ Scoping Review
Table I: Comparison of Technology in Use for Privacy Preservation in IoT
①: The study covers both technical and non-technical PPTs approach
②: The study covers only the technical privacy preservation technique approach.
③: The study introduces new topics and emerging technologies
: The study reviews the topic comprehensively,      $\star$: The study covers the topic partially,
: The study does not cover the topic.
The comparative insights shown in Table I highlight the uniqueness of this scoping review. This study is the first to conduct a comprehensive scoping review of PPT in an IoT system, providing a detailed understanding of the technologies used in building PETs, IoT applications, privacy goals, the implementation of PPT across different layers of IoT computing, and the types of privacy associated with these techniques. We present an explanation of similar reviews related to this study as follows:
Kinza et al. [18] studied how privacy can be protected in IoT applications. They placed a particular emphasis on how fog computing can strengthen privacy in the IoT field. They discussed the PETs used for implementing privacy solutions: anonymization and data access control. Their study primarily focused on technical approaches such as homomorphic encryption, logistical scrambling, and distributive role-based access controls [18]. Additionally, the study clearly explained the types of privacy protection covered by the techniques they examined. While they reviewed many IoT applications, the specific privacy goals of the studied techniques received relatively limited attention.
Abi Sen et al. [19] partially discuss some technologies of use in their study, primarily focusing on mathematical models and cryptography used for privacy preservation in IoT systems in the literature. However, their study covers common privacy goals and IoT applications (except newly introduced IoT applications in multimedia, social networks, and wearables). They primarily emphasized the preservation of user and location privacy as the main focus of their review, highlighting the challenges associated with these specific privacy types in their study [19]. Additionally, they explain the significance of central cloud storage in IoT and its inherent limitations, proposing the adoption of Fog computing to address these limitations through decentralized storage and processing capabilities [19].
Torre et al. [20] addressed a few specific privacy goals of PPTs in their study. These goals include anonymization, access control, data flow privacy, data minimization, and obfuscation. While the authors [20] provide some insights into the technology used for building PETs, their primary focus centers on cryptography in this category. However, they also touch upon blockchain and machine learning as pertinent approaches. Furthermore, their review predominantly focuses on IoT applications in health, transport, and wearable [20] without covering the recently introduced IoT applications.
In their systematic review, Ali et al. [21] focus on privacy-preserving identifiers in IoT environments, particularly those that enable pseudonymity. The primary focus of the study lies on X.509 certificates for cryptography and privacy identifiers that facilitate pseudonym usage and data minimization. Their study of IoT applications is domain-specific, with a specific focus on areas such as health, smart homes, transportation, and energy production [21].
Some gaps were observed in previous reviews on PPTs in the IoT environment, especially in the technology used to build PETs, privacy goals, privacy preservation types, the IoT application of interest, and the PPT implementation in the IoT computing layer. Prior reviews on PPT in IoT focus on studies that use technical approaches to achieve PPT. However, PPTs in IoT should also incorporate non-technical approaches such as informed consent.
Furthermore, this study discovered that no prior reviews have covered new trends in IoT applications in areas such as social networks and multimedia.
This study addressed some of these research gaps by discussing potential solutions (see Section X) based on specific research questions. This study also explored the goals of various privacy techniques, including hybrid solutions that combine multiple approaches. Additionally, this study focused on different types of privacy, especially those that have received less attention from researchers, such as device identity privacy, and analyzed IoT application areas, particularly those recently introduced in IoT.
IV The Five-step Research Methodology
This study adopted PRISMA-ScR (Preferred Reporting Items for Systematic Reviews and Meta-Analyses Extension for Scoping Reviews) due to its comprehensive search strategy, which helps map the literature into meaningful categories within a broad context. The scoping review reporting guideline was initially introduced by Tricco et al. [8]. This guideline features a detailed checklist and explanations of PRISMA-ScR. This study implements PRISMA-ScR within the framework established by Arksey et al. [39]. Arksey’s framework consists of five stages: (1) identifying research questions, (2) identifying relevant studies, (3) selecting pertinent studies, (4) charting the collected data (data visualization), and (5) synthesizing and presenting the findings. It is important to note that this guideline inherently includes the essential components outlined in the PRISMA reporting methodology.
V Step 1: Identifying the research questions
To better understand the factors affecting privacy and its preservation in IoT systems, we formulated essential questions not addressed in previous studies.
These inquiries uncover what, why, and how certain IoT activities and applications influence the techniques employed to preserve privacy in IoT systems. Here, the study presents a set of research questions (RQs) that are not adequately addressed in prior reviews on PPTs in the IoT.
  • RQ1: What technologies are prominently used to develop PETs in IoT systems?
  • RQ2: What are the dominant PPT’s privacy goals in IoT applications?
  • RQ3: What are the commonly known and recent IoT applications added to the literature?
  • RQ4: What computing layer in IoT architectures are PPTs mostly implemented within IoT systems, and which IoT application has PPT most implemented from the literature?
  • RQ5: What privacy types are the main focus of most PPTs in IoT systems?
VI Step 2: Identifying relevant studies
To understand privacy preservation in the IoT environment, this study queried selected databases with the following search expression: [(”Privacy preservation techniques” OR ”Privacy protection techniques” OR ”Privacy preservation methods” OR ”Privacy protection Approaches” OR ”Privacy protection methods” OR ”Privacy protection techniques”) AND (”Internet of Things” OR ”IoT” OR ”Smart Things” OR ”Smart environment” OR ”Internet of Everything”)].
This study relied on reputable sources from seven databases: IEEE Explorer, ACM Digital Library, ScienceDirect, SpringerLink, Sage Journal, IET, and Taylor and Francis, ensuring the robustness and reliability of our findings. As illustrated in Figure 2, a total of 725 papers were initially collected, including conference proceedings, journal articles, white papers, websites, and theses. The first step of this study involved reviewing the titles, during which 223 papers were excluded for reasons such as being white papers, surveys, non-academic documents, or irrelevant content. Next, the abstracts and conclusions of the remaining 502 papers were examined, excluding 245 papers that were not related to the Internet of Things (IoT). From the remaining 257 papers, twelve (12) papers that lacked privacy considerations were excluded, and two duplicates were removed. This process resulted in a final total of 243 papers for our review.
This study applied both inclusion and exclusion criteria. The inclusion criteria required papers to focus on PPT in IoT, propose at least one PPT for IoT, be peer-reviewed, and be published between 2010 and 2023. However, the exclusion criteria ruled out secondary manuscripts, non-IoT-related papers, non-educational papers, white papers, non-scientific documents, and non-English papers.
VII Step 3: Selection of PPT publications in IoT
This study explored seven databases to examine their content on PPTs in IoT systems. IEEE Explorer was the most prolific, contributing about 59% of the total publications. In contrast, Taylor and Francis, Sage and IET had the lowest contributions, as shown in Figures LABEL:fig:Journal-Database-bar and LABEL:fig:Journal-Database-pie.
The volume of PPT publications declined from 2013 to 2017, reached its peak in 2019, and has been increasing again from 2019 to 2023, with another notable peak occurring in 2021. By the first quarter of 2023, which marked the end of our paper collection, 42 published papers had been gathered that year, indicating a significant rise in PPT publications, as illustrated in Figure 4.
Figure 5 shows that IEEE Explorer houses the highest number of PPT papers based on PPT goals in IoT systems, followed by ScienceDirect. Taylor and Francis, IET, and Sage contributed modestly, each offering between one and two PPT papers in IoT.
Figure 2: The flow diagram showing the paper selection criteria
Figure 4: Number of publications on PPTs in IoT per year between 2010-2023
Figure 5: Papers on PPT goals in IoT system and their published Database
VIII Step 4: Charting the data collected on PPT publications
The collected papers are categorized by PPT goals, technology used for building PET, IoT applications utilizing PPTs, computing layers where PPT is integrated within the IoT architecture, and type of privacy protection. Our results will address our research questions in subsection IX by reviewing each category.
VIII-A PPT and PPT Goals
Many studies on IoT privacy have introduced various techniques aimed at different privacy goals. To organize this extensive literature, publications were categorized from various databases by the specific privacy protection goals they addressed.
VIII-A1 Anonymity
Anonymity in data privacy involves altering the original data format to hide sensitive parts of the that could lead to the identification of the data subject [40, 41, 42]. Various techniques are used to achieve anonymity:
  • Data perturbation: Adding random noise to data to preserve its usefulness for analysis while protecting its sensitive content [43].
  • Data Generalization: Reducing data precision to a level that prevents easy linkage to individuals. It is used in maintaining data accuracy for tasks like medical data analysis [44, 45].
  • Data Masking: Concealing original data through modified content, retaining the data format and utility while protecting sensitive information [46, 47].
  • Data Swapping: Exchanging data attributes value with values from different records to preserve statistical characteristics while obscuring individual data records[48] [49].
  • Pseudonymization: Replacing identifiable information with pseudonyms or codes to protect sensitive data, making it difficult to link transformed data back to original identities [50] [51].
VIII-A2 Data flow privacy
This protects sensitive data during transmission processes such as data offloading, sharing, routing, and collaboration.
  • Data Sharing protection: Data sharing is essential in IoT systems but poses privacy risks. Different PPTs are employed to protect sensitive data during transmission and sharing processes [52].
  • Multipath routing: IoT networks like mobile ad hoc networks (MANETs) use multipath routing protocols to enhance privacy protection for location, identity, data, and traffic as proposed by Taheri et al. [53]
  • Data verification protection: This process reviews and helps improve existing privacy implementation in an IoT system. Techniques such as Petri net analysis proposed by Malek et al. [54] are used to analyze and improve privacy implementations in IoT systems, identifying and addressing privacy violations.
  • Route path protection: This is crucial in location-based services in IoT networks. For example, obfuscation of path is proposed by Yang et al. [55] to protect users’ path privacy.
  • Collaborative Privacy protection: Collaboration in distributed architecture is essential. However, trust needs to be established among the entities involved. Therefore, privacy techniques that collectively protect data sharing during collaboration are important to establish trust among the collaborative entities in the IoT environment [56].
VIII-A3 Data Access Control
Access control in IoT systems involves allocating and granting access privileges based on defined policies and models. Ragothaman et al. [57] discuss eight access control models, with seven being the main focus in PPT papers as they are associated with IoT systems:
  • Role-Based Access Control (RBAC): Users are granted access based on their roles within the system, ensuring appropriate access privileges. Managing RBAC in large and dynamic IoT environments can be challenging [57][58].
  • Attribute-Based Access Control (ABAC): The access privilege is based on attributes or features of an IoT device, such as device ID, location, and IP address, to mention a few, and can be used for evaluating the device before access is granted [57] [58].
  • Blockchain-Based Access Control (BBAC): Access control based on blockchain technology includes transaction-based and smart contract-based mechanisms. Transaction-based control allows granting, revoking, or delegating access, while smart contracts enforce access policies [57].
  • Capability-Based Access Control (CapBAC): This approach requires the user or device to possess certain capabilities before the access token is delivered. This approach does not rely on the context before providing access [58].
  • Usage-Based Access Control (UCON): UCON integrates authorization, obligation, and condition concepts. Authorization checks user eligibility, obligation specifies conditions for access, and conditions define criteria that users must fulfill [57] [58].
  • Relationship-based Access Control (ReBAC): ReBAC grants access based on established relationships in IoT systems, such as user-to-user or device-to-device relationships. This approach supports dynamic Identity Relationship Management (IRM) solutions [57].
  • Discretionary Access Control (DAC): DAC grants access based on an Access-Control List (ACL) managed by administrators. It requires careful monitoring and updating, especially when adding or removing devices or users from the IoT network [57][58].
VIII-A4 Data Awareness
Ensuring data subjects are informed and have control over their data, aligning with General Data Protection Regulation (GDPR) policies on user-informed consent [59]. For example, Tamani et al. [60] proposed using a semantic firewall infrastructure to filter personal data access in IoT.
VIII-A5 Data Isolation
It is the process of confining IoT devices or services away from other network segments. Thapa et al. [61] describe different levels of isolation, and Zhang et al. [62] proposed virtualization and containerization in smart healthcare systems.
VIII-A6 Data Minimization
Reduces data collection and processing in IoT systems to ensure only relevant data are used, preventing over-utilization. Studies focus on data minimization in IoT include [63, 64, 65].
VIII-A7 Multiple Goals
Some PPTs aim to achieve multiple goals simultaneously: Data anonymization and access control in [66, 67, 68], data flow privacy and access control in [69], data flow privacy and minimization in [70], data isolation and distortion in [27]. Figure II presents other papers that achieved multiple goals in their proposed work.
Our observations: From our study, anonymity emerges as the primary goal of PPTs across IoT application systems, with data masking being the most commonly used method to achieve this. This observation is supported by our findings as shown in Table III. Additionally, many PPTs address multiple privacy goals simultaneously, as shown in Table II. These findings underscore the importance of achieving anonymity in protecting sensitive data across diverse IoT applications while highlighting the need for comprehensive privacy strategies that simultaneously address multiple facets of data protection.
Privacy Goals Paper
Anonymity and Access control [66] [67] [68] [71] [72]
Data flow protection and Data minimization [70]
Data minimization and Data Isolation [73]
Anonymity and Data isolation [27] [74] [75] [76] [77] [78]
Data flow protection and Access control [69]
Access control and Data Isolation [79]
Anonymity and Data flow protection [80] [81]
Access control and Data Minimization [82] [83]
Table II: Papers on PPT with multiple privacy goals
{tblr}
width = colspec = Q[120]Q[120]Q[320]Q[70], cell21 = r=5, cell71 = r=5, cell121 = r=7, vlines, hline1-2,7,12,20-23 = -, hline3-6,8-11,13-19 = 2-4, PPT Goals & PPT method Paper Number of Paper
Anonymity Data masking [84] [85] [86] [87] [88] [89] [90] [91] [92] [93] [94] [95] [96] [97] [98] [99] [100] [101] [102] [103] [104] [105] [106] [107] [108] [109] [110] [111] [112] [113] [114] [115] [116] [117] [118] [119] [120] [121] [122] [123] [124] [125] [126] [127] [128] [129] [130] [131] [132]   49
Data generalization [133] [134] [135] [136] [137] [138] [139] [140] [141] [142] [143] [144] [145] [146] [147] [148] [149] 17
Data perturbation  [32] [150], [151] [152] [153] [154] [155] [156] [157] [158] [159] [160] [161] [162] [163][164] [165] [166] [167] [168] [25] [169] [170] [171] [172] [24] [173] [174] [175] [176] [177] [178] [179] [180] [181] [182] [183] [184] 38
Data Pseudonymization [185] [186] [187] [188] [189] [190] [191] [192] [193] [194] [195] [23] [196] [197] [198] [199] [200] [201] [202] [203] 20
Data Swapping [204] 1
Data Flow Privacy Data Sharing protection [205] [28] [206] [207] [208] [209] [210] [211] [212] [213] [214] [215] [216] 13
Multi-path routing [217] [218] [219] 3
Data Privacy Verification [220] [221] 2
Route path protection [222] [223] [224] [225] [226] [227] [228] [229] [230] [231] 10
Collaborative PP [232] [233] [234] [235] 4
Access Control RBAC [236] [237] [238] [239] 4
ABAC [240] [241] [242] [26] [243] [244] [245] [246] [247] [248] [249] 11
RBAC and ABAC [250] 1
BBAC [251] [252] [253] [22] [254] [255][256] [257] [258] [259] [260] [261] [262] [263] [264] [265] [266] [267] 18
CapBAC [268] [269] [270] 3
UCON [271] [272] [273] 3
ReBAC [274] 1
DAC [275] [276] [277] 3
Data Privacy Awareness - [278] [279] [280] [281] [282] 5
Data Isolation - [283] [284] [285] [286] [287] 4
Data Minimization - [288] [289] [290] [291] [292] [293] [294] [295] [296] [297] [298] [299] 12
Table III: Papers on PPT Goals and their Approaches
VIII-B Technology used in building PETs
PETs are tools used to create PPTs in a system. These PETs are built using innovative technologies such as blockchain, machine learning/deep learning, and post-quantum algorithms, to mention a few, as shown in Figure 9. Some of these technologies are explained as follows:
VIII-B1 Blockchain
Blockchain technology decentralizes the control system, fostering a distributed and peer-to-peer transaction environment among anonymous parties within a digital ledger system [300] [301]. It enhances privacy in IoT systems through smart contracts, which enable trust transactions that prevent identity disclosure and tracking [302].
Blockchain has emerged as a novel protective layer for the internet and its applications. Privacy is achieved through the provision of authentication, validation, and tracing mechanisms, serving as robust safeguards for assets within the IoT system. [300].
VIII-B2 Cryptography
This approach encodes the data to be accessed by only authorized users. An example is Homomorphic protection, which allows encrypted data to be computed at its encrypted state without decrypting it [229] [234].
VIII-B3 Machine and Deep Learning (MD/DL)
The introduction of AI has allowed the implementation of ML/DL in IoT systems for privacy protection. However, their implementation circle around protecting data in process at the nodes or the cloud (service provider end), as seen in federated learning(FL) applications in IoT privacy protection [298, 290, 284, 235, 212, 140, 23, 118, 87, 151, 27, 24].
VIII-B4 Mathematical Model
Some studies proposed novel mathematical models to protect privacy in the IoT environment. Most of these models center on protecting data in transit during communication among IoT system components. The proposed mathematical model either provides cryptography or access control to the system [32, 270, 187, 97, 100, 220].
VIII-B5 Optimization
This method provides enhanced privacy protection in IoT systems by optimizing existing approaches like lightweight encryption methods suitable for IoT environments [222, 91, 173, 198, 224, 175, 225, 172, 139].
VIII-B6 Post-Quantum Algorithm
This approach utilizes a post-quantum algorithm to ensure quantum-safe privacy protection in IoT. Liu et al. [133] proposed a lattice-based encryption scheme to ensure the encrypted data are quantum-safe. This is the only paper in this study that proposed post-quantum encryption for privacy protection.
VIII-B7 Infrastructure-based
These PPTs leverage the infrastructure component in an IoT system [236, 186, 94, 135, 288, 259, 261]. Some active devices that provide access control are deployed physically or virtually as a Network function (NF) to control who, when, and how subject data is accessed in an IoT system.
VIII-B8 Informed Consent Management
It is the the only non-technological approach in our study. It relies on policies and laws governing data privacy, such as GDPR, Personal Information Protection and Electronic Documents Act (PIPEDA), California Consumer Privacy Act (CCPA), and Brazilian General Data Protection Act (LGPD), as outlined by regulatory bodies [278, 271, 295].
Our observations: Some studies combine some of these PET building technologies to achieve multiple privacy preservation goals. Examples are Singh et al. [140], Fang et al. [118], and Zhao et al. [151] all use ML and blockchain to provide privacy preservation in IoT. Also, Salim et al. [27], and Hindistan [25] both utilize ML and DP to build PET in IoT systems.
VIII-C IoT Applications and PPTs
The literature typically focuses on common IoT applications in transportation, home/building, city, health, and industry. However, IoT applications are expanding into social networks and multimedia, as seen in social IoT and the Internet of Multimedia Things (IoMT).
VIII-C1 Social IoT (S-IoT)
This can be seen in the mutually coordinated social cooperation that exists among IoT devices, with their connection strength depending on interaction levels. This connections are built on trust and privacy [219][186] [218]. These interactions can leverage platforms like Web 3.0, as seen in social media applications such as SM-IoT [27] [296].
VIII-C2 Internet of Multimedia of Things (IoMT)
This is the smart operation in multimedia systems. Privacy protection in the Internet of Multimedia of Things (IoMT) has recently garnered significant attention from researchers. Various methods have been introduced to protect sensitive data. For example, Elmisery et al. [123] proposed a data mashup service tool for aggregating data from web applications, and Gati et al. [297] proposed a deep private tensor train autoencoder (dPTTAE) technique for handling large volumes of data in smart environments. Additionally, Liang et al. [285] proposed a Secure Service Discovery (SSD) approach to protect data communication during the service discovery process, addressing vulnerabilities to eavesdropping and data tampering. Deep learning has emerged as the dominant building technology for PET in IoMT applications, as shown in Figure 6.
Figure 6: Papers on PPT implementation in IoT applications and Technology use to build PETs from 2010 to the first quarter of 2023
IoT Application Edge Fog Cloud Local Edge and Cloud Fog and Cloud Fog, Edge and Cloud Local, Edge, Fog and Cloud Local and Cloud Not Mentioned
Smart Home [151], [264], [265], [131], [83], [72] [270], [148] [278],[236], [223], [224],[288], [71] - [66] - - - - [269], [246], [105], [157], [276], [115]
Industrial IoT (IIoT) [212] [26] [252], [229], [25] [133],[yin2017location] [70] - [156] - - [251], [87], [211]
Smart Grid [165],[215] - [158], [237], [203] - [213] - - - - [289], [92], [160]
Medical IoT [234],[146], [79], [147],[74], [149], [78] [94], [238],[73] [116], [140], [84], [240], [245], [111], [68],[129],[235],[166],[101],[195],[258],[259],[303],[239],[117],[249] [191] [80] - - - - [142],[170],[214],[114]
Social IoT (S-IoT) [219] - [186], [183] - - - - - - [296],[27],[218]
IoT of Vehicle (IoV) [222] [185],[242],[89],[176],[199] [90],[206],[253],[188],[280],[273],[260],[107] [180] [241], [81], [267] [75] - - - [28],[189],[201], [192],[102],[291],[193],[150],[263]
Wearable IoT - - [190], [293], [178] [196] - - - - - [97],[298]
Smart City [113],[272],[210] [204], [182] [32],[86],[283] - [217] - - - [173],[254] [208], [152],[187]
Internet of Multimedia of things (IoMT) - [285] [123] - - - - - - [297]
IoT in Agriculture - - [125] - - - - - - -
IoT in Education - - [167] - - - - - - -
IoT in Ecommerce - - - - - - - - - [99]
General IoT [169],[91],[230],[279],[118],[244],[110],[163],[154],[209],[247],[177],[139],[198], [216], [180], [181], [184], [82], [299], [76], [287], [132], [77] Fog [205],[243],[153], [161],[164],[295],[23],[281], [286], [277] [141],[250],[88],[268],[119],[126],[202],[badsha2018designing],[162],[112],[98],[225],[256],[232],[233],[103],[292],[220],[221],[109] [207],[290],[271],[69], [231], [282] [93],[284],[127] [257] - [24] [227] [85],[122],[143][22],[134],[226],[95],[96], [135],[255],[274],[136],[130],[137],[175],[120],[228],[144],[155],[275],[100],[104],[106],[128],[124],[145],[294],[168],[138],[197],[171],[172],[67],[200],[108],[261],[262],[159],[121], [266], [179], [77]
Table IV: Paper on PPTs and Layer of Implementation in IoT architecture from 2010 to the first quarter of 2023.
Privacy Type Number of papers papers
User Data Privacy 164 [32] [92] [143] [22] [113] [119] [151] [237] [160] [187] [280] [227] [269] [189] [94] [243] [201] [126] [153] [202] [223] [264] [190] [161] [114] [245] [271] [96] [135] [246] [255] [162] [123] [110] [130] [174] [137] [111] [224] [191] [98] [209] [120] [288] [238] [283] [257] [99] [144] [273] [204] [192] [275] [100] [101] [102] [291] [284] [105] [66] [193] [167] [194] [157] [146] [258] [145] [259] [68] [129] [293] [294] [276] [260] [168] [25] [107] [220] [138] [197] [235] [221] [170] [296] [211] [242] [198] [169] [199] [152] [159] [73] [118] [297] [27] [172] [213] [125] [212] [67] [239] [298] [262] [109] [214] [69] [263] [166] [156] [88] [268] [230] [24] [279] [118] [122] [278] [26] [156] [70] [158] [116] [140] [84] [240] [142] [117] [219] [241] [90] [272] [205] [250] [85] [141] [121] [265] [71] [215] [72] [216] [181] [184] [131] [178] [79] [147] [80] [182] [81] [183] [148] [267] [82] [282] [249] [299] [76] [179] [149] [83] [132] [77] [78]
Device Data Privacy 13 [290] [134] [93] [274] [164] [228] [106] [171] [91] [207] [115] [133] [87]
User and Device Data Privacy 24 [208] [173] [163] [97] [112] [233] [103] [104] [295] [139] [23] [281] [270] [200] [108] [261] [236] [251] [252] [289] [234] [186] [74] [203]
Location Privacy 20 [226] [188] [95] [175] [225] [176] [195] [124] [177] [196] [285] [155] [150] [218] [28] [89] [180] [231]
General Privacy 22 [244] [136] [154] [165] [256] [232] [210] [127] [292] [128] [247] [229] [222] [185] [206] [253] [86] [217]
Table V: Publications on privacy protection types in IoT systems from 2010 to the first quarter of 2023.
Our observations: According to Figure 6, the number of papers on PET building technology specific to IoT applications in agriculture, education, and e-commerce is the lowest. In contrast, the fields of Smart Homes, Medical IoT, and Industrial IoT (IIoT) contain a significant number of papers within our study, as illustrated in Figure 6. Additionally, some papers propose PET building technologies that could be applied across all IoT applications. However, this general approach may not be suitable for specific IoT applications, given the differences in architectures, protocols, resource requirements, and communication processes, as noted by Sana et al. [304].
VIII-D PPT implementation in IoT Application and System Architecture
PPT implementation in IoT has been proposed across different IoT architectures. These architectures have different layers where PPTs can be deployed. This study focus on the implementation of PPTs in IoT computing layers as shown in Figure 7.
PPT can be locally deployed within the IoT system at the aggregate (gateway) or at the edge close to the IoT system (which is common in Fifth-Generation (5G) enabled IoT systems). Fog and cloud-based architecture are also commonly used today, and most privacy studies have used some of these layers in IoT architecture to deploy PPTs.
Our observations: Some proposed studies on PPTs are deployed across two or more layers of the IoT architectures. Table IV shows the categorization of papers on PPTs and IoT computing layers in which they are deployed. Research work on end-to-end PPT deployment in IoT architecture that cuts across local, fog, edge, and cloud is limited, as shown in Table IV. Also, it was observed that the cloud is the most used platform where PPTs have been deployed in IoT architecture, followed by edge, then fog. However, many papers did not mention which computing layer they deployed PPT in their publications, as shown in Figure 7.
Figure 7: Publications on PPTs implementation at different computing layers in IoT architecture from 2010 to the first quarter of 2023
VIII-E Privacy Types provided by PPTs in IoT systems
As gathered from this study, the three main privacy types focused on in most papers are user privacy, device privacy, and location privacy. User data privacy protects sensitive data such as Personally Identifiable Information (PII), Protected Health Information (PHI), or any sensitive information that can reveal the identity of a user. The protection of data stored, processed, or transmitted across devices that could reveal either user or device identity in IoT is called device data privacy. Location privacy protects IoT user protection against privacy threats like localization and tracking. This threat also applies to mobile IoT devices like wearable or stationed devices.
Apart from these three privacy types common in IoT systems, some studies have proposed PPTs that simultaneously mitigate user and device privacy. Meanwhile, some studies have presented the general privacy of the entire IoT system without any specific privacy type.
Our observations: This study presents different privacy types proposed from the literature on PPT implementation in IoT systems. User data privacy has the highest number of papers focusing on user data protection. Few papers proposed PPT to protect device data and location privacy, as shown in Table V and Figure 8. This answers our research question RQ6.
Our study observed that device data privacy proposed by most studies still focuses on user data being stored, processed, or transmitted across IoT devices. This data could comprise both user and device data. Therefore, to protect device identity, sensitive device data such as IP address, MAC address, Operating system, and serial number, to mention a few that could reveal device identity, need to be protected. These sensitive data associated with IoT devices could be linked with other data to identify the device in the IoT system.
Figure 8: Papers on Privacy protection types in IoT system from 2010 to the first quarter of 2023
IX Step 5: Collating, Summarizing, and Reporting Results that Address Research Questions
Using the PRISMA-ScR guidelines for selecting relevant papers, our analysis reveals a steady increase in publications on PPTs from 2013 to 2019, with 2019 having the highest number of publications, as shown in Figure 4. There was a noticeable decline in 2020, probably due to the global pandemic. However, publication rates rebounded and continued to rise from 2021 through the first quarter of 2023, when the paper collection was concluded. The following subsections present results that address the research questions outlined in section V.
IX-A RQ1: What technologies are prominently used to develop
PETs in IoT systems? To address RQ1, this study analyzed seven technologies used to build Privacy-Enhancing Technologies (PETs)—Blockchain, Cryptography, Machine and Deep Learning, Mathematical Models, Optimization, Infrastructure-Based techniques, and post-quantum algorithms. Informed consent management is the only non-technical approach to privacy preservation, as gathered from our study. It has also been revealed that cryptography is the most prevalent technology used for privacy preservation in IoT applications, followed closely by blockchain. As observed in this study, the post-quantum algorithm is primarily applied to Industrial IoT (IIoT), indicating its pioneering role and potential for future research in IIoT privacy.
IX-B RQ2: What are the dominant PPT’s privacy goals in IoT applications?
Our study’s research question (RQ2) investigates the dominant goals of PPTs across IoT applications. Our analysis identifies various PPT goals, including anonymity, data flow privacy, access control, privacy awareness, data isolation, and data minimization, as shown in Table III. Among these goals, anonymity, particularly through methods like data masking and data perturbation, emerges as the most prevalent goal in the literature, with a significant percentage of publications focusing on these techniques.
IX-C RQ3: What are the commonly known and recent IoT applications added to the literature?
IoT applications commonly span cities, transportation, health, and agriculture. However, new application areas like Social IoT(S-IoT) and Internet Multimedia of Things (IoMT) are gaining traction as shown in Figure 6. These new platforms directly interact with users and are particularly vulnerable to privacy threats due to their extensive internet exposure.
IX-D RQ4: What computing layer in IoT architectures are PPTs mostly implemented within IoT systems, and which IoT application has PPT most implemented from the literature?
IoT applications have increased recently, with medical IoT having the highest PPTs implementation as seen in Figure 6. This shows the importance of privacy in medical IoT systems. This level of importance is a result of sensitive medical data transactions involved in the MIoT system containing data such as Personally Identifiable Information (PII) and Protected Health Information (PHI). In case of a data breach, these sensitive medical data can reveal the identity of an individual in relation to their medical condition. Furthermore, our observation shows that PPTs are more implemented around the cloud layer in the IoT architecture in most IoT applications compared to other computing layers, as shown in Table IV.
IX-E RQ5: What privacy types are the main focus of most PPTs in IoT systems?
According to our review, the three main privacy types in IoT are user, device, and location privacy. Some studies proposed that PPTs can protect the privacy of both the user and the device, as seen in Table V. However, our observation shows that most privacy implementations in IoT devices mainly focus on protecting user data stored in these IoT devices, with very limited or no emphasis on the importance of preserving the identity of the IoT device that stores these user data. This observation shows that user privacy is the ultimate goal of many research studies on PPT in IoT systems.
IX-F Relationship between Key concepts in IoT system privacy
In this scoping review, we proposed a privacy preservation taxonomy that includes major privacy entities, as explained in Section II. This taxonomy encompasses Privacy Enhancing Technologies (PET), Privacy Protection Techniques (PPT), Privacy Threats, Privacy Goals, and the technologies used to develop PET. Although previous studies have mentioned these entities, none have explored the correlations among them in the context of privacy preservation, particularly within an IoT environment. Figure 9 illustrates the relationships among these privacy entities and detailed examples for each entity.
Figure 9: Privacy terms and their relation based on privacy preservation in IoT
Note: Existing and emerging technologies are utilized to build PET, which is essential for developing PPTs that achieve specific privacy goals in IoT. These privacy goals provide different privacy types. The overall process provides privacy preservation that mitigates privacy threats in IoT systems.
X Future Direction
After proper reviews of PPTs proposed in the selected papers from top literature databases, this study observed some future research areas that need researchers’ attention.
X-A Need for device privacy preservation (Device Identity protection)
Finn et al. [11] and Liyanage et al. [12] have proposed several types of privacy. However, in a typical user service, provider interactions within an IoT system use two major types of information flow: (a) user data, such as personally identifiable information (PII), including contacts, names, photos, and age, and (b) IoT device data and activity, such as IP address, geolocation, and device ID. These two major data are essential; however, protecting the identity of the device is crucial, as its protection could directly impact the privacy of the user and the location [186].
Research Gap Despite limited studies on device data privacy, most studies are centered around protecting user data stored, processed, or transmitted across devices in IoT. Therefore, research on device identity privacy is essential for protecting sensitive information about devices in IoT systems.
X-B Need for proper understanding of privacy threats activities
Providing PPT to mitigate privacy threats in IoT is important, but more important is understanding the activities of these privacy threats. The knowledge of privacy threat activities gives more insight to provide proper PPT for mitigating these privacy threats.
Research Gap Implementing a comprehensive PTA for IoT systems is essential to address the full spectrum of privacy threats. In addition, this study observed a lack of privacy-based datasets to train and test known privacy threats in the Internet of Things. As an interim solution, generating synthetic datasets could simulate typical threat behaviors in IoT systems, helping to build and validate effective privacy protection measures.
X-C Importance of Privacy Engineering (PE), Privacy Threat Analysis (PTA), and Privacy Impact Assessment (PIA) in privacy preservation
PE is the continuous and overall implementation of privacy protection across any IoT system. PE is required in IoT systems to implement privacy protection beyond the design stage. PTA analyzes privacy threats, and PIA assesses the risk level of these privacy threats in IoT systems. These three privacy approaches (PE, PTA, and PIA) are essential in any IoT system. However, an integrated approach that incorporates these privacy tools is essential in selecting appropriate PPTs to mitigate privacy threats in IoT systems.
Research Gap PE, PTA, and PIA integrated approach is important at an early stage of implementing privacy preservation in IoT applications. However, researchers have not effectively emphasized such approach in implementing privacy protection in a wide range of IoT systems. Alalade et al. [305] implemented this integrated approach in the Smart Home (SH) environment. The end result of their work ranked privacy threats based on their level of impact in an SH, enabling suggestions of possible PPTs to mitigate some of these threats. Applying this and similar approaches to other IoT applications should enable proper and in-depth privacy protection.
X-D The influence of emerging technology
The introduction of emerging technologies has a dual impact on IoT systems. On the one hand, technology plays a crucial role in improving privacy protection [6, 7]. On the other hand, it also provides malicious actors with advanced tools to carry out privacy threats [4, 5]. This duality highlights the complex balance between utilizing technological advancements for privacy protection and managing the increased risks presented by the sophisticated tools employed by malicious individuals.
Research Gap Researchers should leverage emerging technology to enhance privacy protection in IoT systems. For instance, integrating an appropriate lightweight post-quantum cryptography algorithm with the assistance of generative AI will provide quantum-safe privacy in an IoT system.
XI Conclusion
This scoping review utilizes PRISMA-ScR guidelines and Arksey et al.’s framework to review PPTs in IoT systems comprehensively.
Out of 725 papers collected, 243 met the selection criteria. These selected papers were analyzed for insights into PPT goals, PET building technologies, PPT implementation in IoT applications, where these PPTs are implemented in IoT computing architectures, and protected privacy types.
Our review discovered that cryptography is the predominant technology for preserving privacy in IoT systems. At the same time, anonymity remains the primary objective of PPT in IoT implementations, addressing our research questions RQ1 and RQ2, respectively. Furthermore, the continuous expansion of IoT applications across diverse domains, with new introductions in social networks as SIoT and multimedia networks as (IoMT) reflects the growing influence of IoT, which answers our RQ3 research question. This study also reveals that PPT is implemented across various IoT computing layers: cloud, fog, edge, and local network. However, most PPT implementation occurs at the cloud layer of the IoT system. In addition, medical IoT has more publications on the implementation of PPT than other IoT applications, as gathered in our study, addressing our research question RQ4. This study discovered that while most PPT implementations prioritize user privacy, fewer studies focus on device and location privacy, thus answering RQ5.
Finally, implementing effective PPT to address privacy threats in IoT systems can be challenging due to IoT system resource constraints commonly associated with these systems. Nevertheless, this study highlights our findings through a comprehensive scoping review, outlines future research directions in PPT implementation in IoT systems, and acknowledges limitations in this work. This review aims to serve as a valuable foundation for researchers in selecting appropriate PPTs for IoT applications and provide an understanding of the privacy types these PPTs protect.
Limitations : The study focused on articles published from 2010 to early 2023, potentially missing relevant studies outside this timeframe. The inclusion criteria limited articles to English-language publications, which may have excluded significant non-English research. Additionally, as a scoping review, the study provided a broad overview rather than an in-depth analysis, potentially omitting essential nuances. These limitations should be considered when interpreting the findings.